# Digital requirements engineering with an INCOSE-derived SysML meta-model

### v1

@@ 6 Conclusions
This paper presented an extension of prior MBSR research using the latest INCOSE Guide to Writing Requirements and relates it an emerging DRE paradigm. Over 300 requirements were written and revised with project stakeholder feedback to test the effectiveness of this approach in improving quality and connectedness contributing to a valuable ASoT. The experience gained through this NASA JPL system development project was shared in brief while presenting a minimal example in the space flight domain. The benefits and challenges listed in subsections 3.1 and 3.2 are gathered through experience, and while the benefits are likely transferable to other tools, the challenges may be ameliorated by improved software performance and enhanced model-based workflows for DRE activities.
@@
By encoding the primary object classes (Fig. 1) from the INCOSE GtWR into the systems architecture modeling tool, rapid improvement in real-world requirements quality was achieved by a systems engineering student intern over a period of 6 months. The INCOSE-derived SysML meta-model embedded in this MBSR Profile supports a DE approach to requirements engineering and system development that may reduce costs and improve quality if deployed at scale. SysML-based architecture modeling tools may facilitate the MBSR approach by supporting requirements management needs such as automatic timestamps, organization-defined collaborative workflows, extensible and rigorous requirement identifier definition, text values with embeddable model elements, and enhanced speed of the software. The authors believe that MBSRs have the potential to leverage SysML strengths in the transition to DE while effectively satisfying stakeholder needs, but further testing, feedback, and experimentation are needed to further validate this claim.
@@ Acknowledgements
The first author would like to thank Paulo J. Younse for his mentorship.
@@
Access the open-source SysML Profiles and example MBSR models on GitHub (The Authors, 2024).

### v2

@@ 6.3 Summary
This paper presented an extension of prior MBSR research using the latest INCOSE Guide to Writing Requirements and related it to an emerging DRE paradigm. Over 300 requirements were written and revised with project stakeholder feedback to test the effectiveness of this approach in improving quality and connectedness, contributing to a valuable ASoT. The experience gained through this NASA JPL system development project was shared in brief in Sec. 4. The benefits and challenges listed in subsections 4.1 and 4.2 were gathered through experience, and while the benefits are likely transferable to other tools, the challenges may be ameliorated by improved software performance and enhanced model-based workflows for DRE activities.
@@
We found that development of SysML requirements in this direction has been fruitful for gaining support of MBSE in the context of requirements engineering. By encoding the primary object classes (Fig. 4) from the INCOSE GtWR into the systems architecture modeling tool, rapid improvement in real-world requirements quality was achieved by a systems engineering student intern over a period of 6 months. The INCOSE-derived MBSR profile supports a DE approach to requirements engineering and system development that may reduce costs and improve quality if deployed at scale. Our MBSR technique is independent of a methodology, as its intent is to extend the usefulness of SysML models that include requirements, from early-phase requirements development to architecture solution development to later-phase SE activities. The technique is therefore adaptable to new and existing SE methodologies. SysML-based architecture modeling tools may facilitate the MBSR approach by supporting requirements management needs such as automatic timestamps, organization-defined collaborative workflows, extensible and rigorous requirement identifier definition, text values with embeddable model elements, and enhanced speed of the software. The authors believe that INCOSE-derived MBSR has the potential to leverage SysML strengths in the transition to DE while effectively satisfying stakeholder needs, but further development, testing, feedback, and experimentation are needed to further validate this claim.
@@ Acknowledgements
The first author would like to thank Paulo J. Younse for his mentorship, and the Mars Returned Sample Handling team for their participation and support.
@@
Access the open-source SysML profiles and example MBSR models on GitHub (herber2024repo).
Appendix A Appendix